2. Kapoor Denims Inc. manufactures denim fabric. 20 metres


of the fabric is sold at Rs 8000. Fixed costs is Rs 20 lacs per


production period and the profit contribution is 40 percent of


price. (20 marks)


a.


Determine the BEP.


b.


Determine the Profit/loss at output of 8,000, 10,000


and 15,000 units.


C.


For the next production period, fixed costs will


increase to 30 lacs due to a major capital investment


programme, but the new and more efficient machinery will


result in a lower variable production cost so that variable cost


per metre will be reduced by 40 percent. If price is


unchanged, re-compute the profit /loss at output rate of 8,000,


10,000 and 15,000 units

1
Expert's answer
2021-12-02T20:14:01-0500

(a)

"BEP=\frac{Fixed Cost}{Contribution Margin}"

"=\frac{20}{0.40\times 8000}=0.00625"

(b)

Profit = Total Revenue- Total Cost

8,000:

"\frac{8000}{20}=400"

"=(400\times8000)-2000000"

"=1200000"


10,000:

"\frac{10000}{20}=500"

"=(500\times8000)-2000000"

"=2000000"


15,000:

"\frac{15000}{20}=750"

"=(750\times8000)-2000000"

"=4000000"


(c)

8,000:

"=(400\times8000)-3000000"

"=200000"


10,000:

"=(500\times8000)-3000000"

"=1000000"


15,000:

"=(750\times8000)-3000000"

"=3000000"